OpenStudy (anonymous):

how do I simplify: 7 square root of 3 end root minus 4 square root of 6 end root plus square root of 48 end root minus square root of 54

OpenStudy (jdoe0001):

\(\large \bf 7\sqrt{3}-4\sqrt{6}+\sqrt{48}-\sqrt{54}?\)

OpenStudy (fibonaccichick666):

So here is my question, can you make every separate piece contain a \(\sqrt{3}\) somehow?